OpenStudy (anonymous):

if u= (1+i square root of 3) and v= (1-i square root of 3). What is uv? A) 1 B) -4 C) -4i D)4

OpenStudy (dan815):

hey

OpenStudy (dan815):

expand the brackets out

OpenStudy (dan815):

you get \[(1)^2+i\sqrt3-i\sqrt3-(i\sqrt3)^2 \] \[(1)^2-i^2(\sqrt3)^2 = 1+3=4 \]